OpenStudy (anonymous):

Factor using the GCF: 2x3 - 4x

OpenStudy (anonymous):

the GCF of the 2 is 2x, so factor 2x out to get 2x(x^2 - 4) maybe go further to say 2x(x+2)(x-2)

OpenStudy (anonymous):

\[2x^3-4x=2 \cdot x \cdot x^2-2 \cdot x \cdot 2\]

OpenStudy (anonymous):

What factors do the terms have in common?

OpenStudy (anonymous):

yea i think so im not sure im new at this stuff

OpenStudy (anonymous):

u seem new

OpenStudy (anonymous):

Find the GCF of all the terms in the algebraic expression. 2. Write the GCF outside the brackets. 3. Divide the terms in the original algebraic expression by the GCF and write them inside the brackets.